HOME ESR EXPERIMENTS LABS etc
E121 E127 E132 E125 E143 laser_cooling_2021 E142 Ê128 E146 E0052 E0018 E0028 E0038 E0075 E0203_E0018
  masses & isomeres 2025  Not logged in ELOG logo
Message ID: 34     Entry time: Tue Apr 1 10:34:20 2025
Author: Jelena B, Siddharth 
Category: General 
Subject: Resolved issue with SIS18 
Beam is back with some fluctuations.
We did not have beam from 9:34 until 10:08, around 35 minutes. 
Attachment 1: 2025-04-01_10-33-38-948_tcl1064.png  425 kB  Uploaded Tue Apr 1 11:35:12 2025  | Hide | Hide all
2025-04-01_10-33-38-948_tcl1064.png
ELOG V3.1.5-fc6679b